OpenStudy (anonymous):

Sketch the right triangle and find the length of the side not given. If necessary, approximate the length to the nearest thousandth. leg = 6, leg = 8

OpenStudy (aroub):

(x)^2=(6)^2+(8)^2 x^2=36+64 x^2=100 x=10